THE MICROWAVE SPECTRUM AND STRUCTURE OF VINYLSELENOL

\underline{G. WLODARCZAK},D. PETITPREZ, J. DEMAISON, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR 8523 CNRS, Université des Sciences et Technologies de Lille, F-59655 Villeneuve d’Ascq, Cedex, France ; H. RIAGUE, J. C. GUILLEMIN, Laboratoire de Synthèse et Activation de Biomolécules, UMR 6052, CNRS, ENSCR, F-35700 Rennes, France.

The microwave spectrum vinylselenol, H_{2}C=CHSeH, has been investigated using a Fourier transform microwave spectrometer in the 6 - 18 GHz region. The ground state of the parent species have been assigned for the two most stable conformers: the syn and the gauche forms. The ground state of several Se and ^{13}C isotopomers were also observed in natural abundance, and deuterated species in enriched samples. For the most abundant species the millimeter wave spectra were also analyzed, leading to a centrifugal distortion analysis. Some structural parameters were determined (including C-Se bond length) and compared to ab initio calculations.